OpenStudy (anonymous):

Select the verbal phrase that corresponds with the given expression. (a + b)3 a plus b Triple the sum of a and b The sum of a and b cubed Three times a plus three times b

OpenStudy (anonymous):

a plus b = a+b sum of a and b cubed = (a+b)^3 three times a plus three times b = (3a+3b) = ... triple the sum of a and b = ...